### v1

# IndiTag: An Online Media Bias Analysis and Annotation System Using Fine-Grained Bias Indicators

Abstract
In the age of information overload and polarized discourse, understanding media bias has become imperative for informed decision-making and fostering a balanced public discourse. This paper presents IndiTag, an innovative online media bias analysis and annotation system that leverages fine-grained bias indicators to dissect and annotate bias in digital content. IndiTag offers a novel approach by incorporating large language models, bias indicator, vector database to automatically detect and interpret bias. Complemented by a user-friendly interface facilitating both automated bias analysis and manual annotation, IndiTag offers a comprehensive platform for in-depth bias examination. We demonstrate the efficacy and versatility of IndiTag through experiments on four datasets encompassing news articles from diverse platforms. Furthermore, we discuss potential applications of IndiTag in fostering media literacy, facilitating fact-checking initiatives, and enhancing the transparency and accountability of digital media platforms. IndiTag stands as a valuable tool in the pursuit of fostering a more informed, discerning, and inclusive public discourse in the digital age. The demonstration video can be accessed from https://youtu.be/Gt2T4T7DYqs. We release an online system for end users and the source code is available at https://github.com/lylin0/IndiTag.
IndiTag: An Online Media Bias Analysis and Annotation System Using Fine-Grained Bias Indicators
[…]
1 Introduction
In an era dominated by online media, understanding and mitigating bias within digital content Yu et al. (2008); Iyyer et al. (2014); Liu et al. (2022a) have become paramount. The rapid dissemination of information through digital platforms has not only facilitated access to a wealth of knowledge but has also underscored the importance of critically evaluating the biases inherent in the content we consume. In response to these challenges, our work introduces IndiTag, a sophisticated system designed to empower users with tools for analyzing and annotating bias in online media content.
At the heart of IndiTag lies a commitment to providing users with a comprehensive and accessible analysis of bias within digital articles. Leveraging advanced techniques from large language models (LLMs), vector database Peng et al. (2023) and IndiVec technique Lin et al. (2024), our system employs a two-stage approach to facilitate bias analysis. In the offline stage, IndiTag constructs a bias indicator vector database using LLMs and meticulously designed prompts. These prompts serve as guides for LLMs in generating fine-grained bias indicators, which are then rigorously verified to ensure quality and reliability.
Moving to the online stage, users interact with IndiTag’s intuitive interface, which presents them with visualizations and explanations of bias within the input article. By visualizing descriptors alongside matched indicators and their corresponding matching scores, users gain immediate insights into the presence and extent of bias within the text. Furthermore, the system highlights the relationship between descriptors and the original text input at the sentence level, allowing users to pinpoint instances of bias and understand their implications within the broader context of the article.
The visual presentation of bias analysis within IndiTag serves to enhance user comprehension by distilling complex information into digestible visual cues. Through clear and intuitive visualizations, users can easily identify and interpret instances of bias, regardless of their level of expertise in bias analysis.
Furthermore, to demonstrate the effectiveness of IndiTag in predicting bias labels, we conduct experiments utilizing various datasets (i.e., FlipBias Chen et al. (2018), BASIL Fan et al. (2019), BABE Spinde et al. (2022), MFC Card et al. (2015)). These experiments serve to validate the meaningfulness of the matched indicators in analyzing bias and highlight the system’s efficacy in accurately predicting bias across diverse datasets.
In brief, by providing users with powerful tools for bias analysis and annotation, IndiTag represents a significant advancement in promoting transparency and accountability in digital media discourse. Our work underscores the importance of equipping users with the means to critically evaluate the biases present in online media content, empowering them to make informed assessments and contribute to a more nuanced understanding of digital information.
2 Related Work
Media bias is a complex and multifaceted phenomenon, encompassing various forms of partiality and skewed perspectives in the presentation of information Golbeck et al. (2017). While definitions and scopes of media bias vary across studies, researchers have approached this topic from diverse angles, including political biases Liu et al. (2022a), linguistic biases Spinde et al. (2022), text-level context biases Färber et al. (2020), gender biases Grosz and Conde-Cespedes (2020), and racial biases Barikeri et al. (2021). Regardless of the specific focus, methodologies typically involve classification approaches, ranging from classical methods like Naive Bayes and Support Vector Machines Evans et al. (2007); Yu et al. (2008); Sapiro-Gheiler (2019) to more advanced techniques such as Recurrent Neural Networks Iyyer et al. (2014) and pretrained language model-based methods like BERT and RoBERTa Liu et al. (2022a); Fan et al. (2019).
Recent research has explored diverse aspects of media bias, including defining it Nakov et al. (2021); Spinde (2021), detecting bias in content Liu et al. (2022b); Lee et al. (2022); Fagni and Cresci (2022); Spinde et al. (2021), understanding bias perception Goldman and Mutz (2011); Knobloch-Westerwick and Jingbo Meng (2009); Barnidge et al. (2020); Gearhart et al. (2020), and promoting bias awareness through educational interventions Weeks et al. (2017); Correll et al. (2004); Kim (2019); Pennycook and Rand (2021). Additionally, studies have explored connections between media bias and related concepts such as misinformation, hate speech, sentiment, and political ideology, while also addressing ethical considerations. This body of literature demonstrates a growing interest to understanding and addressing bias in media content.
3 IndiTag System
3.1 Functionality Overview
Figure 1: Functionality interpretation of IndiTag Online Analysis and Annotation System.
The IndiTag system is designed to empower users with a comprehensive suite of tools for analyzing and annotating bias in online media content. At its core, the system offers automated bias detection using advanced large language models and fine-grained bias indicators, alongside a user-friendly interface for manual annotation and validation. Users can upload digital media content and access detailed visualizations and explanations regarding bias within the content. Fig. 1 illustrates the key functionalities offered by the IndiTag system, including:
  • Support both Single Article and Structured Document Input. Users can input either single articles or upload structured documents (i.e., a JSON file containing articles to be analyzed per line). This input mechanism serves as the starting point for bias analysis within the IndiTag system, facilitating an efficient and user-friendly experience.
  • Automated Bias Analysis. IndiTag conducts automated analysis of the input article, leveraging a database of bias indicators and sophisticated algorithms. Through this process, the system systematically examines the content to identify indicators suggestive of bias. By generating a list of top-k bias indicators, users gain clear insights into perceived bias within the content.
  • Explainable Bias Matching Mechanism. To enhance user comprehension, IndiTag incorporates an explainable bias matching mechanism. Users can directly access or refer to corresponding segments from the original article for each generated bias descriptor. This feature allows pinpointing specific areas contributing to identified bias indicators, facilitating a deeper understanding of bias manifestation within the content.
  • Support for Bias Annotation and Results Downloads: In addition to automated analysis, IndiTag offers robust support for bias annotation. Users can upload documents and engage in online annotation, leveraging the system’s analysis of descriptors and indicators to effectively annotate bias within the content. Following annotation, the system compiles annotated articles, including detailed matched descriptors, indicators, the relationship between descriptors and segments, along with comments from annotators, for convenient downloading.
3.2 Workflow Overview
Figure 2: Interpretation of detailed modules of IndiTag Online Analysis and Annotation System.
As illustrated in Fig. 2, our IndiTag system comprises two stages: offline construction of a bias indicator vector database and online bias analysis and annotation. We briefly introduce these stages below to provide readers with a rough understanding of how our system operates.
Offline Bias Indicator Database.
The vector database is constructed offline to serve as the core module supporting online indicator matching, consisting of sentence embeddings representing bias indicators. Leveraging LLMs, we employ effective prompts to automatically extract bias indicators from a dataset containing bias and corresponding labels (e.g., political left-leaning or right-leaning). Our offline database construction methodology closely follows our previous research paper Lin et al. (2024), with additional engineering efforts (e.g., multi-strategy indicator filtering) and advanced representative indicator selection method (i.e., strict clustering based indicator selection) aimed at enhancing the quality of the indicator set to improve overall bias analysis performance. Following this extraction process, we store the embeddings of the entire indicator set in the database, ready for use in online bias analysis queries.
Online Bias Analysis
The online analysis stage begins by taking the user’s input article and extracting descriptors based on LLMs. These descriptors’ embeddings are then utilized to match against the stored indicators. Our IndiTag system visualizes the descriptors, matched indicators along with their matching scores, and the relationship between descriptors and the original text input at the sentence level. By visualizing the descriptors alongside matched indicators and their corresponding matching scores, users can quickly grasp the presence and extent of bias within the text. Moreover, by displaying the relationship between descriptors and the original text input at the sentence level, users gain insights into where and how bias manifests throughout the article.
3.3 Implementation Details
In this subsection, we delve into the detailed implementation of IndiTag, providing insights into how each component operates to fulfill its functions.
System Input and Output.
The IndiTag system is designed to receive digital media content as input and provide users with comprehensive bias analysis results as output. Users can input articles or structured documents for analysis, and the system generates visualizations and explanations regarding the presence and nature of bias within the content. The output includes descriptors, matched indicators with matching scores, and the relationship between descriptors and the original text input at the sentence level.
3.3.1 Fine-grained Bias Indicator Construct
The construction of the fine-grained bias indicator dataset involves several steps:
Designing Prompts for Indicator Generation.
We meticulously craft prompts to guide LLMs in generating fine-grained bias indicators (refer to details in our prior research work Lin et al. (2024)). These prompts cover various aspects of media bias assessment, including tone, language, sources, coverage, balance, agenda, framing, and bias in examples and analogies.
Bias Indicator Generation.
LLMs are guided by the prompts to generate bias indicators. The generated indicators form the initial set, denoted as $\mathcal{I}_{0}$, which undergo further verification. A multi-strategy verification process is employed to ensure the quality of the generated indicators. Conflicting indicators are eliminated, and indicators with low confidence are excluded through backward verification using LLMs. In the process of indicator verification, we prompt gpt-3.5-turbo-16k model for the confidence score (a number from 1 to 10) of each indicator. The verified indicator set is denoted as $\mathcal{I}_{v}$, with corresponding bias labels $\mathcal{Y}_{v}$.
Strict Clustering-Based Indicator Improvement.
In addition to the intuitive indicator verification methods discussed earlier, we introduce a rigorous clustering-based approach to further refine the quality of bias indicators generated by LLMs. Specifically, given a set of bias indicators $\mathcal{I}_{v}={i_{v1},i_{v2},...,i_{vn}}$, we utilize Euclidean Distance to cluster the embeddings of these indicators into distinct clusters. The number of clusters is controlled by the hyper-parameter $\alpha$, which dictates the maximum distance for indicators to be clustered together. Subsequently, we select the indicator nearest to the center of each cluster as the representative indicator. This strict clustering process serves two main purposes: 1) it effectively reduces redundancy among indicators, ensuring that only the most distinct and informative indicators are retained, and 2) it balances the distribution of indicators across various clusters, preventing an overabundance of indicators within specific semantic spaces. By employing this method, we enhance the overall quality and diversity of bias indicators, thereby improving the reliability and accuracy of bias analysis within the IndiTag system. The filtered indicator set is denoted as $\mathcal{I}$.
3.3.2 Bias Descriptor and Indicator Matching
To facilitate bias analysis, we construct a vector database $\mathcal{V}_{\mathcal{I}}$ based on the indicator set $\mathcal{I}$, storing vector representations of each indicator. These vectors capture the semantic information of the indicators, enabling efficient comparison with query text inputs. Specifically, each indicator $i_{j}$ undergoes embedding extraction to obtain its vector representation $v_{j}$, utilizing techniques such as OpenAI Embeddings.
Given a query text input denoted as $c$, we generate its descriptors $\mathcal{D}^{c}={d_{1}^{c},d_{2}^{c},...,d_{|\mathcal{D}^{c}|}^{c}}$ using the same embedding extraction method. Subsequently, we compute the cosine similarity between each descriptor $d_{j}^{c}$ and the vectors in the indicator database $\mathcal{V}_{\mathcal{I}}$. This process yields a distance metric $Distance(v_{j}^{c},v_{k})$ for each descriptor-indicator pair, quantifying their semantic similarity.
For each descriptor $d_{j}^{c}$, we rank the bias indicators in $\mathcal{I}$ based on their distances, selecting the top $M$ indicators as candidates for bias analysis. The parameter $M$ serves as a hyper-parameter, controlling the number of indicators considered for each descriptor.
3.3.3 Indicator Enhanced Bias Prediction
These selected indicators collectively contribute to predicting bias label for the input text. By employing majority voting among the bias labels associated with the selected indicators, we determine the bias label assigned to the query text input $c$. This approach ensures robust and comprehensive bias analysis, leveraging the semantic information encoded within both descriptors and bias indicators.
4 Experiments
Base Models
FlipBias
BASIL
FT-BERT FT-GPT GPT-ZS GPT-FS IndiVec IndiTag FT-BERT FT-GPT GPT-ZS GPT-FS IndiVec IndiTag
Scores on Biased Instances
Precision 83.6 88.7 63.9 59.9 62.7 62.9 19.1 20.0 39.3 22.4 32.2 35.2
Recall 98.6 93.6 22.1 61.4 71.6 72.4 100 94.9 2.3 44.7 34.9 34.2
F1 90.5 91.1 32.9 60.6 66.9 67.3 32.0 33.0 4.4 29.5 33.5 34.7
Scores on Both Biased and Non-Biased Instances
Micro F1 87.5 90.0 45.8 52.1 57.2 57.6 16.1 25.0 80.7 59.7 73.7 75.2
Macro F1 89.9 89.8 43.7 49.8 53.2 53.4 19.1 23.9 46.8 50.5 58.6 59.7
Base Models
BABE
MFC
FT-BERT FT-GPT GPT-ZS GPT-FS IndiVec IndiTag FT-BERT FT-GPT GPT-ZS GPT-FS IndiVec IndiTag
Scores on Biased Instances
Precision 49.2 37.7 81.9 53.7 62.9 66.4 86.3 85.8 86.5 86.4 86.9 86.4
Recall 99.8 100 20.1 68.6 78.9 69.1 76.4 95.3 37.2 72.9 78.6 80.0
F1 65.9 54.7 32.2 60.2 70.0 67.7 81.1 90.3 52.3 79.1 82.5 83.1
Scores on Both Biased and Non-Biased Instances
Micro F1 49.2 38.0 58.4 55.4 66.7 67.6 69.3 82.5 41.0 66.8 71.4 72.0
Macro F1 33.0 28.2 51.1 54.7 66.3 67.6 50.0 50.3 37.3 49.4 51.8 50.8
Table 1: Comparison results (in %) on four datasets. “FT” means fine-tuning the bias prediction model using the Flipbias training set, followed by reporting the prediction results on the test sets of the four datasets. “ZS” and “FS” refer to zero-shot and few-shot setting, separately.
4.1 Datasets
We primarily conducte experiments on political bias datasets due to their higher visibility and abundance. Our bias indicator vector database’s construction utilizes the FlipBias dataset Chen et al. (2018), sourced from allsides.com in 2018, comprising 2,781 events with articles representing different political leanings.
Additionally, we evaluate our model on three other datasets: BASIL Fan et al. (2019), BABE Spinde et al. (2022), and the Media Frame Corpus (MFC) Card et al. (2015). These datasets are relabeled as Biased and Non-Biased instances following Wessel et al. (2023). FlipBias: Consists of 2,781 events with articles from left, center, and right perspectives, with biased and non-biased labels. BASIL: Contains 100 sets of articles from various sources, annotated for lexical and informational bias at the span level. BABE: Comprises 3,673 sentences from Media Cloud, annotated for bias by expert annotators. MFC: Contains 37,622 articles labeled as “pro”, “neutral”, or “anti”, with “pro” and “anti” tones considered as biased.
4.2 Comparisons
Fine-tuned Language Models.
For the Finetune model, we fine-tune pretrained language models, specifically BERT Devlin et al. (2018) and GPT3.5, using the training set of the FlipBias dataset. We then present the test performance results on the test sets of the four datasets. When conducting the fine-tuning experiments, we fine-tune the model using the pre-trained BERT model Devlin et al. (2018) and the AdamW optimizer Loshchilov and Hutter (2017). This fine-tuning process was facilitated through Hugging Face Wolf et al. (2020), and we specifically employed the BertForSequenceClassification model.
Zero-shot and Few-shot Approaches with ChatGPT.
For the ChatGpt baseline, we employ zero-shot and few-shot approaches to predict bias labels. Input data are directly presented with proper prompts to query ChatGPT for bias prediction.
IndiVec.
IndiVec Lin et al. (2024) represents the state-of-the-art research in political bias prediction tasks. While our IndiTag shares similarities with IndiVec, we distinguish ourselves through the adoption of a rigorous clustering-based approach to refine the quality of generated bias indicators.
4.3 Evaluation and Validation
(a) Main Page for Input
(b) Generated Descriptors
(c) Matched Indicators
(d) Additional Customized Mapping
Figure 3: Key Pages of IndiTag Interface.
The IndiTag system undergoes thorough evaluation to ensure its effectiveness and reliability in analyzing and annotating bias in online media content. Evaluation metrics such as precision, recall, and F1-score are utilized to assess the system’s performance, providing insights into its utility for users. Our evaluation of model performance encompasses two key aspects:
1) Precision, Recall, and F1 Score for Biased Instances: These metrics gauge the system’s ability to detect bias within the dataset, reflecting the effectiveness of generated descriptors and matched indicators in analyzing bias.
2) MicroF1 and MacroF1 for Both Biased and Non-Biased Instances: These metrics provide a comprehensive assessment of the overall prediction capabilities of the models, considering both biased and non-biased instances.
This evaluation framework ensures that the IndiTag system delivers reliable and accurate analysis and annotation of bias in online media content, thereby empowering users to make informed assessments and interpretations.
4.4 Experimental Results
The comparison results of bias prediction are reported in Table 1. Across all datasets, the IndiTag generally outperforms other baselines, showcasing its robustness in bias prediction tasks. Notably, IndiTag achieves competitive scores in terms of precision, recall, and F1 scores, indicating its effectiveness in identifying biased instances while maintaining a balance between precision and recall. Moreover, when considering the Micro F1 and Macro F1 scores for both biased and non-biased instances, the IndiTag demonstrates superior performance compared to other baselines. This suggests that IndiTag not only excels in detecting bias but also maintains high prediction capabilities across diverse instances, showcasing its versatility and reliability in bias prediction tasks.
5 User Interface
We present key pages of our IndiTag interface for users in Fig. 3 for reference. Fig. 3(a) demonstrates that IndiTag accommodates both single article input and structured document uploading. Fig. 3(b) and Fig. 3(c) display the generated descriptors and matched indicators for each descriptor in the input article. The bar beneath each descriptor represents the distribution of probabilities for “left”, “neutral”, and “right” leanings. Additionally, the similarity between each indicator and descriptor is visually represented to aid users in bias analysis. Fig. 3(d) introduces an additional feature enabling users to input personalized descriptors and articles to analyze the article from the user’s perspective (i.e., descriptor).
6 Conclusion
In conclusion, IndiTag represents a significant advancement in the analysis and annotation of bias in online media content. Through a meticulous two-stage approach, our system provides users with a comprehensive and accessible analysis of bias, leveraging advanced techniques from natural language processing and machine learning. By validating the effectiveness of our matched indicators through experiments across diverse datasets, we have demonstrated the meaningfulness of our approach in predicting bias labels. With its intuitive interface and powerful tools for bias analysis, IndiTag empowers users to critically evaluate digital content, contributing to transparency and accountability in digital media discourse.
Limitations
Despite its strengths, IndiTag also has certain limitations that warrant consideration. One limitation is the reliance on pre-existing bias indicators, which may not encompass the full spectrum of biases present in digital content. Additionally, the effectiveness of IndiTag may vary depending on the quality and diversity of the input datasets used for training and validation. Furthermore, while our system provides valuable insights into bias within digital media, it is not immune to errors or biases inherent in the underlying algorithms and models. Finally, IndiTag’s performance may be influenced by factors such as language nuances and cultural contexts, which may limit its applicability across diverse settings.
Ethics Statement
Our work on IndiTag prioritizes ethical considerations throughout the research process. We recognize the importance of responsible and unbiased analysis of digital media content, and as such, we have taken several specific measures to ensure ethical practices. Firstly, to improve the quality of indicators presented to users when they use the system, we design several strategies to filter the indicators and adopt advanced strict clustering-based methods to obtain a representative indicator set. Additionally, we adhere to principles of fairness and inclusivity by actively seeking diverse perspectives and datasets to mitigate biases inherent in our algorithms and ensure equitable representation. Finally, we are committed to ongoing evaluation and refinement of our system through continuous monitoring, user feedback, and independent audits to address ethical concerns and promote responsible usage in the digital media landscape.
Broader Impact Statement
Our work on IndiTag has the potential to have a significant broader impact across various domains. By providing users with powerful tools for analyzing and annotating bias in online media content, our system promotes transparency and accountability in digital media discourse. This has implications for media literacy, enabling individuals to critically evaluate digital content and make informed decisions. Furthermore, IndiTag can empower journalists, researchers, and policymakers to identify and address biases within media narratives, fostering a more nuanced understanding of complex social issues. Additionally, our commitment to inclusivity and fairness ensures that IndiTag can be applied across diverse cultural and linguistic contexts, making it accessible to a global audience. Overall, our work has the potential to contribute to a more informed and equitable digital media landscape, ultimately promoting democratic values and fostering constructive dialogue in society.

### v2

@@
-IndiTag: An Online Media Bias Analysis and Annotation System Using Fine-Grained Bias Indicators
+IndiTag: An Online Media Bias Analysis System Using Fine-Grained Bias Indicators for News Consumers
 Abstract
 In the age of information overload and polarized discourse, understanding media bias has become imperative for informed decision-making and fostering a balanced public discourse.
-This paper presents IndiTag, an innovative online media bias analysis and annotation system that leverages fine-grained bias indicators to dissect and annotate bias in digital content.
-IndiTag offers a novel approach by incorporating large language models, bias indicator, vector database to automatically detect and interpret bias.
-Complemented by a user-friendly interface facilitating both automated bias analysis and manual annotation, IndiTag offers a comprehensive platform for in-depth bias examination.
+However, without the experts’ analysis, it is hard for the readers to distinguish bias from the news articles.
+This paper presents IndiTag, an innovative online media bias analysis system that leverages fine-grained bias indicators to dissect and distinguish bias in digital content.
+IndiTag offers a novel approach by incorporating large language models, bias indicators, and vector database to detect and interpret bias automatically.
+Complemented by a user-friendly interface facilitating automated bias analysis for readers, IndiTag offers a comprehensive platform for in-depth bias examination.
 We demonstrate the efficacy and versatility of IndiTag through experiments on four datasets encompassing news articles from diverse platforms.
@@
 IndiTag stands as a valuable tool in the pursuit of fostering a more informed, discerning, and inclusive public discourse in the digital age.
-The demonstration video can be accessed from https://youtu.be/Gt2T4T7DYqs.
+The demonstration video can be accessed from https://youtu.be/3Tux8CW46OE.
 We release an online system for end users and the source code is available at https://github.com/lylin0/IndiTag.
-IndiTag: An Online Media Bias Analysis and Annotation System Using Fine-Grained Bias Indicators
-[…]
+IndiTag: An Online Media Bias Analysis System Using Fine-Grained Bias Indicators for News Consumers
+Luyang Lin1,2, Lingzhi Wang3thanks:   Lingzhi Wang is the corresponding author., Jinsong Guo4, Jing Li5, Kam-Fai Wong1,2 1The Chinese University of Hong Kong, China 2MoE Key Laboratory of High Confidence Software Technologies, China 3Harbin Institute of Technology, Shenzhen, China 4University College London, UK 5Department of Computing, The Hong Kong Polytechnic University, China 1,2{lylin, kfwong}@se.cuhk.edu.hk, 3wanglingzhi@hit.edu.cn 4jinsong.guo@ucl.ac.uk,5jing-amelia.li@polyu.edu.hk
 1 Introduction
@@
 The rapid dissemination of information through digital platforms has not only facilitated access to a wealth of knowledge but has also underscored the importance of critically evaluating the biases inherent in the content we consume.
-In response to these challenges, our work introduces IndiTag, a sophisticated system designed to empower users with tools for analyzing and annotating bias in online media content.
+In response to these challenges, our work introduces IndiTag, a sophisticated system designed to empower news consumers with tools for analyzing bias in online media content, which only be done by journalism experts in the past.
 At the heart of IndiTag lies a commitment to providing users with a comprehensive and accessible analysis of bias within digital articles.
@@
 These experiments serve to validate the meaningfulness of the matched indicators in analyzing bias and highlight the system’s efficacy in accurately predicting bias across diverse datasets.
-In brief, by providing users with powerful tools for bias analysis and annotation, IndiTag represents a significant advancement in promoting transparency and accountability in digital media discourse.
+In brief, by providing news consumers with powerful tools for bias analysis, IndiTag helps to build the logistical thinking ability of readers and represents a significant advancement in promoting transparency and accountability in digital media discourse.
 Our work underscores the importance of equipping users with the means to critically evaluate the biases present in online media content, empowering them to make informed assessments and contribute to a more nuanced understanding of digital information.
@@
 This body of literature demonstrates a growing interest to understanding and addressing bias in media content.
+Although many experts try to organize and categorize biased articles for the readers, like AllSides 111www.allsides.com, Ad Fontes Media 222adfontesmedia.com, and Media bias / Fact Check 333mediabiasfactcheck.com, however, the analyzed articles are limited and lack of timeliness, which diminishes the role in helping the readers to distinguish the bias.
+To fill in this gap, IndiTag is developed for the news consumers to used directly, so that the readers can analyze the articles they are reading timely.
 3 IndiTag System
 3.1 Functionality Overview
-Figure 1: Functionality interpretation of IndiTag Online Analysis and Annotation System.
-The IndiTag system is designed to empower users with a comprehensive suite of tools for analyzing and annotating bias in online media content.
-At its core, the system offers automated bias detection using advanced large language models and fine-grained bias indicators, alongside a user-friendly interface for manual annotation and validation.
+Figure 1: Functionality interpretation of IndiTag Online Analysis and Note-taking System.
+The IndiTag system is designed to empower users with a comprehensive suite of tools for analyzing and distinguishing bias in online media content.
+At its core, the system offers automated bias detection using advanced large language models and fine-grained bias indicators, alongside a user-friendly interface for news consumers to record and validate the results.
 Users can upload digital media content and access detailed visualizations and explanations regarding bias within the content.
-Fig.
-1 illustrates the key functionalities offered by the IndiTag system, including:
+Fig.˜1 illustrates the key functionalities offered by the IndiTag system, including:
   • Support both Single Article and Structured Document Input.
@@
 This feature allows pinpointing specific areas contributing to identified bias indicators, facilitating a deeper understanding of bias manifestation within the content.
-  • Support for Bias Annotation and Results Downloads: In addition to automated analysis, IndiTag offers robust support for bias annotation.
-Users can upload documents and engage in online annotation, leveraging the system’s analysis of descriptors and indicators to effectively annotate bias within the content.
-Following annotation, the system compiles annotated articles, including detailed matched descriptors, indicators, the relationship between descriptors and segments, along with comments from annotators, for convenient downloading.
+  • Support for Notes-Taking and Results Downloads: In addition to automated analysis, IndiTag offers support for the news consumers to take reading notes.
+Users can upload documents and engage in online bias note-taking, leveraging the system’s analysis of descriptors and indicators to effectively distinguish bias within the content.
+Following adding the notes, the system compiles the analyzed articles, including detailed matched descriptors, indicators, and the relationship between descriptors and segments, along with comments from users, for convenient downloading.
 3.2 Workflow Overview
-Figure 2: Interpretation of detailed modules of IndiTag Online Analysis and Annotation System.
-As illustrated in Fig.
-2, our IndiTag system comprises two stages: offline construction of a bias indicator vector database and online bias analysis and annotation.
+Figure 2: Interpretation of detailed modules of IndiTag Online Analysis and Note-taking System.
+As illustrated in Fig.˜2, our IndiTag system comprises two stages: offline construction of a bias indicator vector database and online bias analysis and note-taking.
 We briefly introduce these stages below to provide readers with a rough understanding of how our system operates.
@@
 This approach ensures robust and comprehensive bias analysis, leveraging the semantic information encoded within both descriptors and bias indicators.
-4 Experiments
+4 System Evaluation
 Base Models
@@
 4.1 Datasets
-We primarily conducte experiments on political bias datasets due to their higher visibility and abundance.
+We primarily conduct the system evaluation on political bias datasets due to their higher visibility and abundance.
 Our bias indicator vector database’s construction utilizes the FlipBias dataset Chen et al.
@@
 Figure 3: Key Pages of IndiTag Interface.
-The IndiTag system undergoes thorough evaluation to ensure its effectiveness and reliability in analyzing and annotating bias in online media content.
+The IndiTag system undergoes thorough evaluation to ensure its effectiveness and reliability in analyzing bias in online media content.
 Evaluation metrics such as precision, recall, and F1-score are utilized to assess the system’s performance, providing insights into its utility for users.
@@
 2) MicroF1 and MacroF1 for Both Biased and Non-Biased Instances: These metrics provide a comprehensive assessment of the overall prediction capabilities of the models, considering both biased and non-biased instances.
-This evaluation framework ensures that the IndiTag system delivers reliable and accurate analysis and annotation of bias in online media content, thereby empowering users to make informed assessments and interpretations.
+This evaluation framework ensures that the IndiTag system delivers reliable and accurate analysis bias aligned with humans in online media content, thereby empowering users to make informed assessments and interpretations.
 4.4 Experimental Results
-The comparison results of bias prediction are reported in Table 1.
+The comparison results of bias prediction are reported in Table˜1.
 Across all datasets, the IndiTag generally outperforms other baselines, showcasing its robustness in bias prediction tasks.
@@
 5 User Interface
-We present key pages of our IndiTag interface for users in Fig.
-3 for reference.
-Fig.
-3(a) demonstrates that IndiTag accommodates both single article input and structured document uploading.
-Fig.
-3(b) and Fig.
-3(c) display the generated descriptors and matched indicators for each descriptor in the input article.
+We present key pages of our IndiTag interface for users in Fig.˜3 for reference.
+Fig.˜3(a) demonstrates that IndiTag accommodates both single article input and structured document uploading.
+Fig.˜3(b) and Fig.˜3(c) display the generated descriptors and matched indicators for each descriptor in the input article.
 The bar beneath each descriptor represents the distribution of probabilities for “left”, “neutral”, and “right” leanings.
 Additionally, the similarity between each indicator and descriptor is visually represented to aid users in bias analysis.
-Fig.
-3(d) introduces an additional feature enabling users to input personalized descriptors and articles to analyze the article from the user’s perspective (i.e., descriptor).
+Fig.˜3(d) introduces an additional feature enabling users to input personalized descriptors and articles to analyze the article from the user’s perspective (i.e., descriptor).
 6 Conclusion
-In conclusion, IndiTag represents a significant advancement in the analysis and annotation of bias in online media content.
+In conclusion, IndiTag represents a significant advancement in analyzing the bias of online media content for news consumers.
 Through a meticulous two-stage approach, our system provides users with a comprehensive and accessible analysis of bias, leveraging advanced techniques from natural language processing and machine learning.
-By validating the effectiveness of our matched indicators through experiments across diverse datasets, we have demonstrated the meaningfulness of our approach in predicting bias labels.
+By validating the effectiveness of our matched indicators through evaluating across diverse datasets, we have demonstrated the meaningfulness of our approach in predicting bias labels.
 With its intuitive interface and powerful tools for bias analysis, IndiTag empowers users to critically evaluate digital content, contributing to transparency and accountability in digital media discourse.
@@
 Our work on IndiTag has the potential to have a significant broader impact across various domains.
-By providing users with powerful tools for analyzing and annotating bias in online media content, our system promotes transparency and accountability in digital media discourse.
+By providing users with powerful tools for analyzing bias in online media content, our system promotes transparency and accountability in digital media discourse.
 This has implications for media literacy, enabling individuals to critically evaluate digital content and make informed decisions.

### v3

@@
-IndiTag: An Online Media Bias Analysis System Using Fine-Grained Bias Indicators for News Consumers
+IndiTag: An Online Media Bias Analysis System Using Fine-Grained Bias Indicators
 Abstract
@@
 IndiTag stands as a valuable tool in the pursuit of fostering a more informed, discerning, and inclusive public discourse in the digital age.
-The demonstration video can be accessed from https://youtu.be/3Tux8CW46OE.
 We release an online system for end users and the source code is available at https://github.com/lylin0/IndiTag.
-IndiTag: An Online Media Bias Analysis System Using Fine-Grained Bias Indicators for News Consumers
-Luyang Lin1,2, Lingzhi Wang3thanks:   Lingzhi Wang is the corresponding author., Jinsong Guo4, Jing Li5, Kam-Fai Wong1,2 1The Chinese University of Hong Kong, China 2MoE Key Laboratory of High Confidence Software Technologies, China 3Harbin Institute of Technology, Shenzhen, China 4University College London, UK 5Department of Computing, The Hong Kong Polytechnic University, China 1,2{lylin, kfwong}@se.cuhk.edu.hk, 3wanglingzhi@hit.edu.cn 4jinsong.guo@ucl.ac.uk,5jing-amelia.li@polyu.edu.hk
-1 Introduction
-In an era dominated by online media, understanding and mitigating bias within digital content Yu et al.
-(2008); Iyyer et al.
-(2014); Liu et al.
-(2022a) have become paramount.
+I Introduction
+In an era dominated by online media, understanding and mitigating bias within digital content [1, 2, 3] have become paramount.
 The rapid dissemination of information through digital platforms has not only facilitated access to a wealth of knowledge but has also underscored the importance of critically evaluating the biases inherent in the content we consume.
-In response to these challenges, our work introduces IndiTag, a sophisticated system designed to empower news consumers with tools for analyzing bias in online media content, which only be done by journalism experts in the past.
+In response to these challenges, our work introduces IndiTag, a sophisticated system designed to empower news consumers with tools for analyzing media bias, specifically political bias, in online media content, which only be done by journalism experts in the past.
 At the heart of IndiTag lies a commitment to providing users with a comprehensive and accessible analysis of bias within digital articles.
-Leveraging advanced techniques from large language models (LLMs), vector database Peng et al.
-(2023) and IndiVec technique Lin et al.
-(2024), our system employs a two-stage approach to facilitate bias analysis.
+Leveraging advanced techniques from large language models (LLMs), vector database [4] and IndiVec technique [5], our system employs a two-stage approach to facilitate bias analysis.
 In the offline stage, IndiTag constructs a bias indicator vector database using LLMs and meticulously designed prompts.
@@
 Furthermore, the system highlights the relationship between descriptors and the original text input at the sentence level, allowing users to pinpoint instances of bias and understand their implications within the broader context of the article.
-The visual presentation of bias analysis within IndiTag serves to enhance user comprehension by distilling complex information into digestible visual cues.
-Through clear and intuitive visualizations, users can easily identify and interpret instances of bias, regardless of their level of expertise in bias analysis.
-Furthermore, to demonstrate the effectiveness of IndiTag in predicting bias labels, we conduct experiments utilizing various datasets (i.e., FlipBias Chen et al.
-(2018), BASIL Fan et al.
-(2019), BABE Spinde et al.
-(2022), MFC Card et al.
-(2015)).
+The visual presentation of bias analysis within IndiTag enhances user comprehension by distilling complex information into digestible visual cues.
+Through clear and intuitive visualizations, users can easily identify and interpret instances of bias, regardless of their expertise level in bias analysis.
+Furthermore, to demonstrate the effectiveness of IndiTag in predicting bias labels, we conduct experiments utilizing various datasets (i.e., FlipBias [6], BASIL [7], BABE [8], MFC [9]).
 These experiments serve to validate the meaningfulness of the matched indicators in analyzing bias and highlight the system’s efficacy in accurately predicting bias across diverse datasets.
 In brief, by providing news consumers with powerful tools for bias analysis, IndiTag helps to build the logistical thinking ability of readers and represents a significant advancement in promoting transparency and accountability in digital media discourse.
-Our work underscores the importance of equipping users with the means to critically evaluate the biases present in online media content, empowering them to make informed assessments and contribute to a more nuanced understanding of digital information.
-2 Related Work
-Media bias is a complex and multifaceted phenomenon, encompassing various forms of partiality and skewed perspectives in the presentation of information Golbeck et al.
-(2017).
-While definitions and scopes of media bias vary across studies, researchers have approached this topic from diverse angles, including political biases Liu et al.
-(2022a), linguistic biases Spinde et al.
-(2022), text-level context biases Färber et al.
-(2020), gender biases Grosz and Conde-Cespedes (2020), and racial biases Barikeri et al.
-(2021).
-Regardless of the specific focus, methodologies typically involve classification approaches, ranging from classical methods like Naive Bayes and Support Vector Machines Evans et al.
-(2007); Yu et al.
-(2008); Sapiro-Gheiler (2019) to more advanced techniques such as Recurrent Neural Networks Iyyer et al.
-(2014) and pretrained language model-based methods like BERT and RoBERTa Liu et al.
-(2022a); Fan et al.
-(2019).
-Recent research has explored diverse aspects of media bias, including defining it Nakov et al.
-(2021); Spinde (2021), detecting bias in content Liu et al.
-(2022b); Lee et al.
-(2022); Fagni and Cresci (2022); Spinde et al.
-(2021), understanding bias perception Goldman and Mutz (2011); Knobloch-Westerwick and Jingbo Meng (2009); Barnidge et al.
-(2020); Gearhart et al.
-(2020), and promoting bias awareness through educational interventions Weeks et al.
-(2017); Correll et al.
-(2004); Kim (2019); Pennycook and Rand (2021).
-Additionally, studies have explored connections between media bias and related concepts such as misinformation, hate speech, sentiment, and political ideology, while also addressing ethical considerations.
+Our work underscores the importance of equipping users with means to critically evaluate biases in online media content, empowering them to make informed assessments and contribute to a more nuanced understanding of digital information.
+II Related Work
+Media bias is a complex and multifaceted phenomenon, encompassing various forms of partiality and skewed perspectives in the presentation of information [10].
+While definitions and scopes of media bias vary across studies, researchers have approached this topic from diverse angles, including political biases [3], linguistic biases [8], text-level context biases [11], gender biases [12], and racial biases [13].
+Regardless of the specific focus, methodologies typically involve classification approaches, ranging from classical methods like Naive Bayes and Support Vector Machines [14, 1, 15] to more advanced techniques such as Recurrent Neural Networks [2] and pretrained language model-based methods like BERT and RoBERTa [3, 7].
+Recent research has explored diverse aspects of media bias, including defining it [16, 17, 18], detecting bias in content [19, 20, 21, 22, 23], understanding bias perception [24, 25, 26, 27], and promoting bias awareness through educational interventions [28, 29, 30, 31].
+Additionally, studies have explored connections between media bias and related concepts such as misinformation, hate speech, sentiment, and political ideology, while also addressing ethical considerations [32].
 This body of literature demonstrates a growing interest to understanding and addressing bias in media content.
 Although many experts try to organize and categorize biased articles for the readers, like AllSides 111www.allsides.com, Ad Fontes Media 222adfontesmedia.com, and Media bias / Fact Check 333mediabiasfactcheck.com, however, the analyzed articles are limited and lack of timeliness, which diminishes the role in helping the readers to distinguish the bias.
-To fill in this gap, IndiTag is developed for the news consumers to used directly, so that the readers can analyze the articles they are reading timely.
-3 IndiTag System
-3.1 Functionality Overview
+To fill in this gap, IndiTag is developed for the news consumers to use directly so that the readers can analyze the articles they are reading timely.
+III IndiTag System
+III-A Functionality Overview
 Figure 1: Functionality interpretation of IndiTag Online Analysis and Note-taking System.
@@
 Users can upload digital media content and access detailed visualizations and explanations regarding bias within the content.
-Fig.˜1 illustrates the key functionalities offered by the IndiTag system, including:
+Fig.
+1 illustrates the key functionalities offered by the IndiTag system, including:
   • Support both Single Article and Structured Document Input.
@@
 By generating a list of top-k bias indicators, users gain clear insights into perceived bias within the content.
+    Figure 2: Interpretation of detailed modules of IndiTag Online Analysis and Note-taking System.
   • Explainable Bias Matching Mechanism.
@@
 This feature allows pinpointing specific areas contributing to identified bias indicators, facilitating a deeper understanding of bias manifestation within the content.
-  • Support for Notes-Taking and Results Downloads: In addition to automated analysis, IndiTag offers support for the news consumers to take reading notes.
+  • Support for Notes-Taking and Results Downloads.
+In addition to automated analysis, IndiTag offers support for the news consumers to take reading notes.
 Users can upload documents and engage in online bias note-taking, leveraging the system’s analysis of descriptors and indicators to effectively distinguish bias within the content.
 Following adding the notes, the system compiles the analyzed articles, including detailed matched descriptors, indicators, and the relationship between descriptors and segments, along with comments from users, for convenient downloading.
-3.2 Workflow Overview
-Figure 2: Interpretation of detailed modules of IndiTag Online Analysis and Note-taking System.
-As illustrated in Fig.˜2, our IndiTag system comprises two stages: offline construction of a bias indicator vector database and online bias analysis and note-taking.
+III-B Workflow Overview
+As illustrated in Fig.
+2, our IndiTag system comprises two stages: offline construction of a bias indicator vector database and online bias analysis and note-taking.
 We briefly introduce these stages below to provide readers with a rough understanding of how our system operates.
@@
 Leveraging LLMs, we employ effective prompts to automatically extract bias indicators from a dataset containing bias and corresponding labels (e.g., political left-leaning or right-leaning).
-Our offline database construction methodology closely follows our previous research paper Lin et al.
-(2024), with additional engineering efforts (e.g., multi-strategy indicator filtering) and advanced representative indicator selection method (i.e., strict clustering based indicator selection) aimed at enhancing the quality of the indicator set to improve overall bias analysis performance.
+Our offline database construction methodology closely follows our previous research paper [5], with additional engineering efforts (e.g., multi-strategy indicator filtering) and advanced representative indicator selection method (i.e., strict clustering based indicator selection) aimed at enhancing the quality of the indicator set to improve overall bias analysis performance.
 Following this extraction process, we store the embeddings of the entire indicator set in the database, ready for use in online bias analysis queries.
-Online Bias Analysis
-The online analysis stage begins by taking the user’s input article and extracting descriptors based on LLMs.
-These descriptors’ embeddings are then utilized to match against the stored indicators.
+  • Online Bias Analysis.
+The online analysis stage begins by taking user’s input article and extracting descriptors based on LLMs.
+These descriptors’ embeddings are then utilized to match against stored indicators.
 Our IndiTag system visualizes the descriptors, matched indicators along with their matching scores, and the relationship between descriptors and the original text input at the sentence level.
-By visualizing the descriptors alongside matched indicators and their corresponding matching scores, users can quickly grasp the presence and extent of bias within the text.
-Moreover, by displaying the relationship between descriptors and the original text input at the sentence level, users gain insights into where and how bias manifests throughout the article.
-3.3 Implementation Details
+By visualizing descriptors alongside matched indicators and their corresponding matching scores, users can quickly grasp the presence and extent of bias within the text.
+Moreover, by displaying the relationship between descriptors and the original text at the sentence level, users gain insights into where and how bias manifests throughout the article.
+III-C Implementation Details
 In this subsection, we delve into the detailed implementation of IndiTag, providing insights into how each component operates to fulfill its functions.
@@
 The output includes descriptors, matched indicators with matching scores, and the relationship between descriptors and the original text input at the sentence level.
-3.3.1 Fine-grained Bias Indicator Construct
-The construction of the fine-grained bias indicator dataset involves several steps:
+  • Fine-grained Bias Indicator Construct. The construction of the fine-grained bias indicator dataset involves several steps:
     Designing Prompts for Indicator Generation.
-We meticulously craft prompts to guide LLMs in generating fine-grained bias indicators (refer to details in our prior research work Lin et al.
-(2024)).
-These prompts cover various aspects of media bias assessment, including tone, language, sources, coverage, balance, agenda, framing, and bias in examples and analogies.
+We meticulously craft prompts to guide LLMs in generating fine-grained bias indicators (see details in our prior work [5]).
+These prompts cover multiple aspects of media bias assessment, including tone and language, sources and citations, coverage and balance, agenda and framing, and bias in examples and analogies.
+The detailed prompt is shown below:
+    Demonstration of bias indicator categories: DESC&EX.
+    Based on the demonstration provided above, please label the TEXT INPUT with bias indicators to identify the political leaning GIVEN LABEL.
+    where DESC&EX represents the description and examples of indicator categories, consistent with those used in [5].
+Building upon the prompt used for generating the indicators, we design a similar prompt for descriptor generation to ensure that the descriptors of the input text share a consistent style with the indicators stored in our database.
+    The definition of bias indicator here is a concise, descriptive label or tag to represent the presence or nature of media bias.
+Human analyze bias indicator from 5 bias indicator categories: Tone and Language, Sources and Citations, Coverage and Balance, Agenda and Framing, Examples and Analogies.
+    Give the TEXT INPUT TEXT
+    Could you summarize the bias indicator from the given five categories and list the bias indicator to identify the political leaning (left, right and Neutral)?
+Here are some output examples EXAMPLES
+    Please reply bias indicator in the following form strictly:
+    caetgory - generated bias indicator (less than 30 words) - political leaning
     Bias Indicator Generation.
@@
     Strict Clustering-Based Indicator Improvement.
-In addition to the intuitive indicator verification methods discussed earlier, we introduce a rigorous clustering-based approach to further refine the quality of bias indicators generated by LLMs.
-Specifically, given a set of bias indicators $\mathcal{I}_{v}={i_{v1},i_{v2},...,i_{vn}}$, we utilize Euclidean Distance to cluster the embeddings of these indicators into distinct clusters.
+In addition to the intuitive indicator verification methods discussed earlier, we introduce a rigorous clustering-based approach [33] to further refine the quality of bias indicators generated by LLMs.
+Specifically, given a set of bias indicators $\mathcal{I}_{v}={i_{v1},i_{v2},...,i_{vn}}$, we utilize Euclidean Distance to cluster the embeddings of these indicators into distinct clusters according to Agglomerative Clustering [34, 35].
 The number of clusters is controlled by the hyper-parameter $\alpha$, which dictates the maximum distance for indicators to be clustered together.
@@
 The filtered indicator set is denoted as $\mathcal{I}$.
-3.3.2 Bias Descriptor and Indicator Matching
-To facilitate bias analysis, we construct a vector database $\mathcal{V}_{\mathcal{I}}$ based on the indicator set $\mathcal{I}$, storing vector representations of each indicator.
-These vectors capture the semantic information of the indicators, enabling efficient comparison with query text inputs.
+  • Bias Descriptor and Indicator Matching.
+    We construct a vector database $\mathcal{V}_{\mathcal{I}}$ based on the indicator set $\mathcal{I}$, storing vector representations of each indicator.
+These vectors capture the semantic information of indicators, enabling efficient comparison with query text inputs.
 Specifically, each indicator $i_{j}$ undergoes embedding extraction to obtain its vector representation $v_{j}$, utilizing techniques such as OpenAI Embeddings.
@@
 The parameter $M$ serves as a hyper-parameter, controlling the number of indicators considered for each descriptor.
-3.3.3 Indicator Enhanced Bias Prediction
+  • Indicator Enhanced Bias Prediction.
 These selected indicators collectively contribute to predicting bias label for the input text.
@@
 This approach ensures robust and comprehensive bias analysis, leveraging the semantic information encoded within both descriptors and bias indicators.
-4 System Evaluation
+  • Additional Customized Mapping.
+To help the user better understand the results through the relationship between the descriptor and the news text, we design a descriptor–text mapping module that visually links descriptors to their occurrences in the original text.
+This module allows users to input both a descriptor and a piece of news text; the system then automatically highlights the segments of the text that correspond to the given descriptor.
+This interactive functionality helps users intuitively examine how specific contextual expressions contribute to media bias.
+The prompt used in this module is shown as below.
+    The descriptors are the key points that may reflect the media bias of an article.
+And the descriptors are considered from 5 bias indicator categories: Tone and Language, Sources and Citations, Coverage and Balance, Agenda and Framing, Examples and Analogies.
+    Give the TEXT INPUT TEXT
+    and a DESCRIPTOR DEP
+    Could you point out the phrases or sentences that reflect this descriptor? Please answer in the format [],[],[][\,],[\,],[\,].
+TABLE I: Comparison results (in %) on four datasets.
+“FT” means fine-tuning the bias prediction model using the Flipbias training set, followed by reporting the prediction results on the test sets of the four datasets.
+“ZS” and “FS” refer to zero-shot and few-shot setting, separately.
 Base Models
@@
 Macro F1 33.0 28.2 51.1 54.7 66.3 67.6 50.0 50.3 37.3 49.4 51.8 50.8
-Table 1: Comparison results (in %) on four datasets.
-“FT” means fine-tuning the bias prediction model using the Flipbias training set, followed by reporting the prediction results on the test sets of the four datasets.
-“ZS” and “FS” refer to zero-shot and few-shot setting, separately.
-4.1 Datasets
+IV System Evaluation
+IV-A Datasets
 We primarily conduct the system evaluation on political bias datasets due to their higher visibility and abundance.
-Our bias indicator vector database’s construction utilizes the FlipBias dataset Chen et al.
-(2018), sourced from allsides.com in 2018, comprising 2,781 events with articles representing different political leanings.
-Additionally, we evaluate our model on three other datasets: BASIL Fan et al.
-(2019), BABE Spinde et al.
-(2022), and the Media Frame Corpus (MFC) Card et al.
-(2015).
-These datasets are relabeled as Biased and Non-Biased instances following Wessel et al.
-(2023).
+Our bias indicator vector database’s construction utilizes the FlipBias dataset [6], sourced from allsides.com in 2018, comprising 2,781 events with articles representing different political leanings.
+Additionally, we evaluate our model on three other datasets: BASIL [7], BABE [8], and the Media Frame Corpus (MFC) [9].
+These datasets are relabeled as Biased and Non-Biased instances following [36].
 FlipBias: Consists of 2,781 events with articles from left, center, and right perspectives, with biased and non-biased labels.
@@
 MFC: Contains 37,622 articles labeled as “pro”, “neutral”, or “anti”, with “pro” and “anti” tones considered as biased.
-4.2 Comparisons
+IV-B Comparisons
   • Fine-tuned Language Models.
-For the Finetune model, we fine-tune pretrained language models, specifically BERT Devlin et al.
-(2018) and GPT3.5, using the training set of the FlipBias dataset.
+For the Finetune model, we fine-tune pretrained language models, specifically BERT [37] and GPT3.5, using the training set of the FlipBias dataset.
 We then present the test performance results on the test sets of the four datasets.
-When conducting the fine-tuning experiments, we fine-tune the model using the pre-trained BERT model Devlin et al.
-(2018) and the AdamW optimizer Loshchilov and Hutter (2017).
-This fine-tuning process was facilitated through Hugging Face Wolf et al.
-(2020), and we specifically employed the BertForSequenceClassification model.
+When conducting the fine-tuning experiments, we fine-tune the model using the pre-trained BERT model [37] and the AdamW optimizer [38].
+This fine-tuning process was facilitated through Hugging Face [39], and we specifically employed the BertForSequenceClassification model.
   • Zero-shot and Few-shot Approaches with ChatGPT.
-For the ChatGpt baseline, we employ zero-shot and few-shot approaches to predict bias labels. Input data are directly presented with proper prompts to query ChatGPT for bias prediction.
+For the ChatGpt baseline, we employ zero-shot and few-shot approaches to predict bias labels.
+Input data are directly presented with proper prompts to query ChatGPT for bias prediction.
   • IndiVec.
-IndiVec Lin et al.
-(2024) represents the state-of-the-art research in political bias prediction tasks.
+IndiVec [5] represents the state-of-the-art research in political bias prediction tasks.
 While our IndiTag shares similarities with IndiVec, we distinguish ourselves through the adoption of a rigorous clustering-based approach to refine the quality of generated bias indicators.
-4.3 Evaluation and Validation
+IV-C Evaluation and Validation
 (a) Main Page for Input
@@
 Figure 3: Key Pages of IndiTag Interface.
-The IndiTag system undergoes thorough evaluation to ensure its effectiveness and reliability in analyzing bias in online media content.
+The IndiTag system undergoes thorough evaluation to ensure its effectiveness and reliability in analyzing media content bias.
 Evaluation metrics such as precision, recall, and F1-score are utilized to assess the system’s performance, providing insights into its utility for users.
@@
 This evaluation framework ensures that the IndiTag system delivers reliable and accurate analysis bias aligned with humans in online media content, thereby empowering users to make informed assessments and interpretations.
-4.4 Experimental Results
-The comparison results of bias prediction are reported in Table˜1.
+IV-D Experimental Results
+The comparison results of bias prediction are reported in Table I.
 Across all datasets, the IndiTag generally outperforms other baselines, showcasing its robustness in bias prediction tasks.
@@
 This suggests that IndiTag not only excels in detecting bias but also maintains high prediction capabilities across diverse instances, showcasing its versatility and reliability in bias prediction tasks.
-5 User Interface
-We present key pages of our IndiTag interface for users in Fig.˜3 for reference.
-Fig.˜3(a) demonstrates that IndiTag accommodates both single article input and structured document uploading.
-Fig.˜3(b) and Fig.˜3(c) display the generated descriptors and matched indicators for each descriptor in the input article.
+It is worth noting that in Table I, we compared several models fine-tuned on the FlipBias dataset.
+These models tend to perform well only on FlipBias itself or on MFC, which shares a similar article-level structure.
+In contrast, although IndiTag is also developed based on FlipBias, it achieves consistently strong performance across all datasets, demonstrating its high adaptability.
+Moreover, IndiTag is constructed without a fine-tuning process, allowing it to be easily updated through minor modifications to the indicator database as media bias patterns evolve over time.
+IV-E Ablation Study
+One of the key improvements from IndiVec to Inditag is the introduction of the strict clustering module.
+The results of the ablation study are presented in Table II.
+As shown, applying our indicator framework leads to a significant increase in recall scores, while the addition of the strict clustering module further improves the F1 scores, resulting in a more stable and reliable system.
+TABLE II: Ablation study results (in %) on four datasets.
+Base Models
+FlipBias
+BASIL
+Pre Rec F1 Pre Rec F1
+Full Model 62.9 72.4 67.3 35.2 34.2 34.7
+   - $\mathcal{I}$ Clustering 62.7 71.6 66.9 32.2 34.9 33.5
+       - $\mathcal{I}$ Indicator 63.9 22.1 32.9 39.3 2.3 4.4
+Base Models
+BABE
+MFC
+Pre Rec F1 Pre Rec F1
+Full Model 66.4 69.1 67.7 86.4 80.0 83.1
+   - $\mathcal{I}$ Clustering 62.9 78.9 70.0 86.9 78.6 82.5
+       - $\mathcal{I}$ Indicator 81.9 20.1 32.2 86.5 37.2 52.3
+V User Interface
+V-A User Interface Design
+We present key pages of our IndiTag interface for users in Fig.
+3 for reference.
+Fig.
+3(a) demonstrates that IndiTag accommodates both single article input and structured document uploading.
+Fig.
+3(b) and Fig.
+3(c) display generated descriptors and matched indicators for each descriptor in the input.
 The bar beneath each descriptor represents the distribution of probabilities for “left”, “neutral”, and “right” leanings.
 Additionally, the similarity between each indicator and descriptor is visually represented to aid users in bias analysis.
-Fig.˜3(d) introduces an additional feature enabling users to input personalized descriptors and articles to analyze the article from the user’s perspective (i.e., descriptor).
-6 Conclusion
+Fig.
+3(d) introduces an additional feature enabling users to input personalized descriptors and articles to analyze the article from the user’s perspective (i.e., descriptor).
+V-B Case Study
+Fig. 3 presents a case illustrating the detailed process of system usage from the perspective of end users. The original news text is as follows:
+A Joe Biden presidency could reset ties with top US trade partner Mexico that have suffered since Donald Trump made his first White House bid, tarring Mexican migrants as rapists and gun runners and vowing to keep them out with a border wall.
+During the descriptor generation phase, the LLM accurately summarizes this passage with statements such as “Describes Trump’s actions negatively and emphasizes negative impact on US-Mexico relations” and “Focuses on negative aspects of Trump’s actions without presenting counterarguments”.
+Subsequently, our system matches these descriptors with the bias indicators in the database, such as “Using negative language to describe Donald Trump’s actions and behavior”.
+Most matched indicators are associated with left-leaning labels, leading the system to assign a left-leaning prediction to the text.
+Furthermore, the interface visually highlights different leanings using distinct colors, allowing users to easily interpret and verify the analysis results.
+This visualization provides an intuitive, user-friendly experience for studying media bias.
+VI Conclusion
 In conclusion, IndiTag represents a significant advancement in analyzing the bias of online media content for news consumers.
@@
 One limitation is the reliance on pre-existing bias indicators, which may not encompass the full spectrum of biases present in digital content.
-Additionally, the effectiveness of IndiTag may vary depending on the quality and diversity of the input datasets used for training and validation.
+Additionally, the effectiveness of IndiTag may vary depending on the quality and diversity of the input datasets.
 Furthermore, while our system provides valuable insights into bias within digital media, it is not immune to errors or biases inherent in the underlying algorithms and models.
+As large language models continue to evolve, further experiments can be conducted to examine whether the model currently used still outperforms other models.
 Finally, IndiTag’s performance may be influenced by factors such as language nuances and cultural contexts, which may limit its applicability across diverse settings.
@@
 Overall, our work has the potential to contribute to a more informed and equitable digital media landscape, ultimately promoting democratic values and fostering constructive dialogue in society.
+Acknowledge
+This work is partially supported by Hong Kong RGC GRF No.
+14206324 and CUHK Knowledge Transfer Project Fund No.
+KPF23GWP20.
+Thanks for the support of the MoE Key Laboratory of High Confidence Software Technologies (MoE Lab).
+We are also grateful to the anonymous reviewers for their comments.

### v4

@@
 I Introduction
-In an era dominated by online media, understanding and mitigating bias within digital content [1, 2, 3] have become paramount.
+In an era dominated by online media, understanding and mitigating bias within digital content [yu2008classifying, iyyer2014political, liu2022politics] have become paramount.
 The rapid dissemination of information through digital platforms has not only facilitated access to a wealth of knowledge but has also underscored the importance of critically evaluating the biases inherent in the content we consume.
@@
 At the heart of IndiTag lies a commitment to providing users with a comprehensive and accessible analysis of bias within digital articles.
-Leveraging advanced techniques from large language models (LLMs), vector database [4] and IndiVec technique [5], our system employs a two-stage approach to facilitate bias analysis.
+Leveraging advanced techniques from large language models (LLMs), vector database [peng2023embedding] and IndiVec technique [lin2024indivec], our system employs a two-stage approach to facilitate bias analysis.
 In the offline stage, IndiTag constructs a bias indicator vector database using LLMs and meticulously designed prompts.
@@
 Through clear and intuitive visualizations, users can easily identify and interpret instances of bias, regardless of their expertise level in bias analysis.
-Furthermore, to demonstrate the effectiveness of IndiTag in predicting bias labels, we conduct experiments utilizing various datasets (i.e., FlipBias [6], BASIL [7], BABE [8], MFC [9]).
+Furthermore, to demonstrate the effectiveness of IndiTag in predicting bias labels, we conduct experiments utilizing various datasets (i.e., FlipBias [chen2018learning], BASIL [fan2019plain], BABE [spinde2022neural], MFC [card2015media]).
 These experiments serve to validate the meaningfulness of the matched indicators in analyzing bias and highlight the system’s efficacy in accurately predicting bias across diverse datasets.
@@
 II Related Work
-Media bias is a complex and multifaceted phenomenon, encompassing various forms of partiality and skewed perspectives in the presentation of information [10].
-While definitions and scopes of media bias vary across studies, researchers have approached this topic from diverse angles, including political biases [3], linguistic biases [8], text-level context biases [11], gender biases [12], and racial biases [13].
-Regardless of the specific focus, methodologies typically involve classification approaches, ranging from classical methods like Naive Bayes and Support Vector Machines [14, 1, 15] to more advanced techniques such as Recurrent Neural Networks [2] and pretrained language model-based methods like BERT and RoBERTa [3, 7].
-Recent research has explored diverse aspects of media bias, including defining it [16, 17, 18], detecting bias in content [19, 20, 21, 22, 23], understanding bias perception [24, 25, 26, 27], and promoting bias awareness through educational interventions [28, 29, 30, 31].
-Additionally, studies have explored connections between media bias and related concepts such as misinformation, hate speech, sentiment, and political ideology, while also addressing ethical considerations [32].
+Media bias is a complex and multifaceted phenomenon, encompassing various forms of partiality and skewed perspectives in the presentation of information [golbeck2017large].
+While definitions and scopes of media bias vary across studies, researchers have approached this topic from diverse angles, including political biases [liu2022politics], linguistic biases [spinde2022neural], text-level context biases [farber2020multidimensional], gender biases [grosz2020automatic], and racial biases [barikeri2021redditbias].
+Regardless of the specific focus, methodologies typically involve classification approaches, ranging from classical methods like Naive Bayes and Support Vector Machines [evans2007recounting, yu2008classifying, sapiro2019examining] to more advanced techniques such as Recurrent Neural Networks [iyyer2014political] and pretrained language model-based methods like BERT and RoBERTa [liu2022politics, fan2019plain].
+Recent research has explored diverse aspects of media bias, including defining it [nakov_survey_2021, spinde_interdisciplinary_2021, spinde2025leveraging], detecting bias in content [liu_politics_2022, lee_neus_2022, fagni_fine-grained_2022, Spinde2021f, lin2023data], understanding bias perception [goldman_friendly_2011, knobloch-westerwick_looking_2009, barnidge_politically_2020, gearhart_hostile_2020], and promoting bias awareness through educational interventions [weeks_incidental_2017, correll_affirmed_2004, kim_hostile_2019, pennycook_psychology_2021].
+Additionally, studies have explored connections between media bias and related concepts such as misinformation, hate speech, sentiment, and political ideology, while also addressing ethical considerations [lin2025counterspeech].
 This body of literature demonstrates a growing interest to understanding and addressing bias in media content.
@@
 Leveraging LLMs, we employ effective prompts to automatically extract bias indicators from a dataset containing bias and corresponding labels (e.g., political left-leaning or right-leaning).
-Our offline database construction methodology closely follows our previous research paper [5], with additional engineering efforts (e.g., multi-strategy indicator filtering) and advanced representative indicator selection method (i.e., strict clustering based indicator selection) aimed at enhancing the quality of the indicator set to improve overall bias analysis performance.
+Our offline database construction methodology closely follows our previous research paper [lin2024indivec], with additional engineering efforts (e.g., multi-strategy indicator filtering) and advanced representative indicator selection method (i.e., strict clustering based indicator selection) aimed at enhancing the quality of the indicator set to improve overall bias analysis performance.
 Following this extraction process, we store the embeddings of the entire indicator set in the database, ready for use in online bias analysis queries.
@@
     Designing Prompts for Indicator Generation.
-We meticulously craft prompts to guide LLMs in generating fine-grained bias indicators (see details in our prior work [5]).
+We meticulously craft prompts to guide LLMs in generating fine-grained bias indicators (see details in our prior work [lin2024indivec]).
 These prompts cover multiple aspects of media bias assessment, including tone and language, sources and citations, coverage and balance, agenda and framing, and bias in examples and analogies.
@@
     Based on the demonstration provided above, please label the TEXT INPUT with bias indicators to identify the political leaning GIVEN LABEL.
-    where DESC&EX represents the description and examples of indicator categories, consistent with those used in [5].
+    where DESC&EX represents the description and examples of indicator categories, consistent with those used in [lin2024indivec].
 Building upon the prompt used for generating the indicators, we design a similar prompt for descriptor generation to ensure that the descriptors of the input text share a consistent style with the indicators stored in our database.
@@
     Strict Clustering-Based Indicator Improvement.
-In addition to the intuitive indicator verification methods discussed earlier, we introduce a rigorous clustering-based approach [33] to further refine the quality of bias indicators generated by LLMs.
-Specifically, given a set of bias indicators $\mathcal{I}_{v}={i_{v1},i_{v2},...,i_{vn}}$, we utilize Euclidean Distance to cluster the embeddings of these indicators into distinct clusters according to Agglomerative Clustering [34, 35].
+In addition to the intuitive indicator verification methods discussed earlier, we introduce a rigorous clustering-based approach [feng2025graph] to further refine the quality of bias indicators generated by LLMs.
+Specifically, given a set of bias indicators $\mathcal{I}_{v}={i_{v1},i_{v2},...,i_{vn}}$, we utilize Euclidean Distance to cluster the embeddings of these indicators into distinct clusters according to Agglomerative Clustering [day1984efficient, feng2024llmedgerefine].
 The number of clusters is controlled by the hyper-parameter $\alpha$, which dictates the maximum distance for indicators to be clustered together.
@@
 We primarily conduct the system evaluation on political bias datasets due to their higher visibility and abundance.
-Our bias indicator vector database’s construction utilizes the FlipBias dataset [6], sourced from allsides.com in 2018, comprising 2,781 events with articles representing different political leanings.
-Additionally, we evaluate our model on three other datasets: BASIL [7], BABE [8], and the Media Frame Corpus (MFC) [9].
-These datasets are relabeled as Biased and Non-Biased instances following [36].
+Our bias indicator vector database’s construction utilizes the FlipBias dataset [chen2018learning], sourced from allsides.com in 2018, comprising 2,781 events with articles representing different political leanings.
+Additionally, we evaluate our model on three other datasets: BASIL [fan2019plain], BABE [spinde2022neural], and the Media Frame Corpus (MFC) [card2015media].
+These datasets are relabeled as Biased and Non-Biased instances following [wessel2023introducing].
 FlipBias: Consists of 2,781 events with articles from left, center, and right perspectives, with biased and non-biased labels.
@@
   • Fine-tuned Language Models.
-For the Finetune model, we fine-tune pretrained language models, specifically BERT [37] and GPT3.5, using the training set of the FlipBias dataset.
+For the Finetune model, we fine-tune pretrained language models, specifically BERT [devlin2018bert] and GPT3.5, using the training set of the FlipBias dataset.
 We then present the test performance results on the test sets of the four datasets.
-When conducting the fine-tuning experiments, we fine-tune the model using the pre-trained BERT model [37] and the AdamW optimizer [38].
-This fine-tuning process was facilitated through Hugging Face [39], and we specifically employed the BertForSequenceClassification model.
+When conducting the fine-tuning experiments, we fine-tune the model using the pre-trained BERT model [devlin2018bert] and the AdamW optimizer [loshchilov2017decoupled].
+This fine-tuning process was facilitated through Hugging Face [wolf2020transformers], and we specifically employed the BertForSequenceClassification model.
   • Zero-shot and Few-shot Approaches with ChatGPT.
@@
   • IndiVec.
-IndiVec [5] represents the state-of-the-art research in political bias prediction tasks.
+IndiVec [lin2024indivec] represents the state-of-the-art research in political bias prediction tasks.
 While our IndiTag shares similarities with IndiVec, we distinguish ourselves through the adoption of a rigorous clustering-based approach to refine the quality of generated bias indicators.